Assume the random variable X has a binomial distribution with the given probability of obtaining a success. Find the following probability, given the number of trials and the probability of obtaining a success. Round your answer to four decimal places. P(X≥7), n=10, p=0.5
Answer #1

Solution

Given that ,

p = 0.5

q = 1 - p = 1 - 0.5 = 0.5

n = 10

Using binomial probability formula ,

P(X = x) = (n C x) * p x * (1 - p)n - x

P(X >7 ) = 1 - P( x <7)

= 1 - P(X = 0) -P(X = 0) -P(X = 0) -P(X = 0) -P(X = 0) -P(X = 0) -P(X = 0)

= 1 - (10 C 0) * (0.5)0 * (0.5)10 - (10 C 1) * (0.5)1 * (0.5)9  - (10 C 2) * (0.5)2 * (0.5)8- (10 C 3) * (0.5)3 * (0.5)7 - (10 C 4) * (0.5)4 * (0.5) 6- (10 C 5) * (0.5)5 * (0.5) - (10 C 6) * (0.5)6 * (0.5)4

=0.1719

  probability=0.1719